ID |
Date |
Author |
Subject |
181
|
Sat Jan 29 11:04:07 2022 |
JM, OH | Grounding and Pulser |
Grounding
Ground cables have been daisy chained together from FEE to FEE. Ground cables are screwed into a feee bolt hole on the FEE body using an M3 bolt.
LEMO cables have been connected in a similar way between the grounds on the adaptor boards for each FEE. Short LEMO cables go between J6 on pairs of adaptor boards in the same FEE holder and long LEMO cables go between J7 on adaptor boards in different FEE holders.
The LEMO grounds from the adaptor boards and the ground cables attached to the FEE bodies are then connected together and then connected to the CRYRING grounding cable.
Pulser
Jan has kindly lent us a pulser. We send the pulser to an inverter giving us a +ve and -ve signal.
The +ve signal goes into the pn FEEs (1,2,5,6)
The -ve signal goes into the nn FEEs (3,4,78)
LEMO cable connecting FEEs in the same holder using J3 on adaptor board. J4 is used for recieving the pulse signal, connecting FEEs in different holders or terminating the signal. |
Attachment 1: Silkscreen_Top.pdf
|
|
180
|
Thu Jan 27 21:33:40 2022 |
JM, OH | AIDA config files, remote access, motor test |
Power, Ethernet and HDMI cables connected to FEEs. To prevent cables coming out of connectors, power and HDMI cables are zip tied to FEE cooling pipe (see attached image).
Correct FEE MAC addresses have been updated in /etc/dhcp/dhcpd.conf
Remote access to carme sever has been setup. Remote access instructions in elogs 176-179.
New ASIC settings which reflect polarity of FEES produced 2022Jan27-16-43-00
New Options file updated to point to new ASIC settings.
/MIDAS/config/TclHttpd/carme-gsi@8015/startup.tcl updated for 8 FEEs
/MIDAS/Linux/startup/NewMerger Updated line from "./master64 -i 16 -l 16 -p 11001 &" to "./master64 -i 8 -l 16 -p 11001 &"
Quick motors test performed. Motors accesible remotely (elog 179). Motors moving output (BNC 1 on motors box) connected to motors module (BNC 1). When not moving +12V is observed at BNC 2 on motors module. When moving 0V is observed on BNC 2. |
Attachment 1: 20220127_101513.jpg
|
|
179
|
Thu Jan 27 15:24:02 2022 |
OH, JM | Access to Motors PI |
1. ssh <user>@lxpool.gsi.de
2. ssh <user or carme>@atppc022.gsi.de
3. ssh pi@CARMEmotorspi (Standard npg password) |
177
|
Thu Jan 27 15:19:18 2022 |
OH, JM | Enabling internet access on the carme-gsi server |
To use both anydesk and firefox on the carme server require an internet connection.
As the server is on the cryring local network an ssh tunnel needs to be opened to the GSI proxy server.
To do this: ssh -L 8080:proxy.gsi.de:8080 carme@atppc025` and leave terminal open and running.
Both anydesk and firefox have been configured to use the proxy localhost:8080
For the carme password ask J. Marsh or C. Bruno |
176
|
Thu Jan 27 14:04:34 2022 |
OH, JM | Remote access to CARME server |
1. ssh into lxpool network: ssh <user>@lxpool.gsi.de
2. ssh into atppc022.gsi.de: ssh carme@atppc022.gsi.de (This is an rpi access server. Don't run code)
3. ssh into carme server: ssh npg@carme-gsi
For the password to the carme user account on the cryring network please contact O. Hall or J. Marsh. |
175
|
Wed Jan 26 19:16:18 2022 |
JM, OH | Water cooling, FEE MAC address |
Water cooling has been connected to the FEE's. Flow meter installed to outlet manifold by way of two right angled pieces. No leaks observed rom any points in the coolant loop, loop will be bled using taps tomorrow. Water turned OFF overnight.
FEE's powered up with the MAC addresses of each read. 1 is the very bottom FEE and 8 the very top. FEE's have also labelled with the MAC addresses.
1. 0xd8 0x80 0x39 0x42 0x0d 0x0c
2. d8:80:39:41:d8:2b
3. d8:80:39:41:d0:05
4. 0xd8 0x80 0x39 0x41 0xd7 0xc2 (Now AIDA 5 see elog 185)
5. 0xd8 0x80 0x39 0x41 0xb4 0x0b (Now AIDA 4 see elog 185)
6. d8:80:39:41:d8:20
7. 0xd8 0x80 0x39 0x41 0xb4 0x16
8 .0xd8 0x80 0x39 0x41 0xd7 0xcd
MACB firmware has been updated by Nic
One ion pump showed the warning 'arcing' (see attached photo) and was subsequently turned OFF. Most likely cause is the cable was moved during installlation of the flow meter as the pump showed no error this morning. The pump was turned back ON, no error observed -> cable not damaged but sensitive to movement. |
Attachment 1: 20220126_162435.jpg
|
|
Attachment 2: 20220126_162448.jpg
|
|
Attachment 3: 20220126-142809.jpg
|
|
174
|
Sat Jan 15 15:38:25 2022 |
JM, OH, TD | FEE installation, cable management |
All 8 FEEs have been installed in their holders.
The damaged feedthrough does not have the D-connector on the adaptor board inserted into the falnge but the FEE is connected to the adaptor board.
FEE's were carefully installed, with one person inserting FEE and another making sure FEE was straight and on the correct pin positions. The spring lock on the bottom FEE's have been tightend and zip ties have been placed to make sure they are secured.
Rack has been populated with cables from the rack to chamber laid out. Cables mostly go above the rack and hang off the large frame. This frame or the CARME frame takes most of the weight of the cables before they are plugged into the FEEs.
Cables are numbered 1-8. Brown ethernet cables are numbered 1-4, green 5,6 and yellow 7,8. Numbers correspnd to Bottom FEE (1,2), Side bottom FEE (3,4), Side top FEE (5,6), Top FEE (7,8)
|
Attachment 1: 20220114_104819.jpg
|
|
Attachment 2: 20220114_105152.jpg
|
|
Attachment 3: 20220114_105213.jpg
|
|
Attachment 4: 20220115_133808.jpg
|
|
Attachment 5: 20220115_133820.jpg
|
|
173
|
Sat Jan 15 11:28:10 2022 |
JM | Inventory |
Vaccum
DN250CF 1x gaskets
DN200CF 5x gaskets
DN160CF 1x gaskets, 2x blank flanges
DN100CF 69x gaskets, 19x blank flanges, 14 double feedthroughs (1 damaged)
DN63CF 3x gaskets, 2x blank flanges
DN300CF 5x gaskets
DN350CF 0x gaskets
1X WIRE SEAL
DN35CF 3x gaskets, 13 blank flanges
DN16CF 22x silver gaskets, 27x copper gaskets
Swagelock
SS-QTM2-D-8M0 Male quick disconnect X18
SS-QTM2A-B-8M0 Female quick disconnect X21
SS-8M0-1-6RS Manifold adaptor X39
NY-8M3-1 Front ferrule X118
NY-8M4-1 Back ferrule X120
Stainless ferrules XMANY
FEE/FEE HOLDER PARTS
FEES x1
Fully assembled FEE holder (MINUS ADAPTOR BOARDS) x1
Cylindrical supports x32
Feedthrough mounts x12
Black adaptor board mounts x11
Spring loaded FEE lock x26
FEE holder side panels x5 pairs
Adaptor boards x2 (1 slightly bent pins)
FEE rail bolts x4
M3X16 281-013 x91
M3 locknuts 260-5815 x34
Springs 821-273 x40
M4x25 (2mm too long) 281-057 x32
M6x20 281-120 x69
M3x6 280-981 x34
Assembled Hose 1.5M (Red) x8
Assembled Hose 1.5M (Blue) x8
Blue Hose ~16M
Red Hose ~10M
We require additional gaskets for those we are low on (<5) DN35,63,160,250,350 and more wire seals
We should have enough swagelock and FEE holder parts to construct another 4 FEE holders for the remaining required for mounting all moving detector FEEs.
We also need fuses for the USB mains relay and power supply -> 2x 250v 10A, 16x 250v 1.6A. Dimensions for both are the same, length=20mm, Diameter=5mm (images attached)
Additional fittings provided by GSI for the manifold are in the swagelock box. |
Attachment 1: 20220115_162649.jpg
|
|
Attachment 2: 20220115_162712.jpg
|
|
172
|
Fri Jan 14 10:00:18 2022 |
TD, OH, JM | Friday 14 January |
11.00 Per https://elog.ph.ed.ac.uk/CARME/170 and https://elog.ph.ed.ac.uk/CARME/171 issues have been identified with the FEE64 carriers and adaptor PCBs
Some additional comments
1) Per https://elog.ph.ed.ac.uk/CARME/170 attachments 6-9 we can see that there are no cable strain reliefs to which we can attach the FEE64 power, RJ45 network and
timestamp HDMI cabling. It would probably be straightforward to design something which could be attached to the FEE64 carrier side panels. For the time being we will use
the copper piping of each FEE64 as the nearest, mechanically stable point to which we secure the cabling.
2) Before the skimming of 1x Delron rail yesterday there was a small amount (<1mm say?) of transverse movement of the FEE64 during insertion into the carrier and it is
not clear why. Small sample of measurements of FEE64 carriers, spacers etc.
Delron adaptor PCB support widths 92.85, 92.87, 92.90
Al DN100CF flange attachment width 92.97, 92.99, 93.01
Al FEE64 carrier spacing rod 93.03, 93.04, 93.03
FEE64 width (w/ 1x c. 2mm skimmed Delron rail) 100.2, 100.2, 100.3 (3 measurements along the length of the 1x FEE64)
FEE64 carrier width 102.8, 103.0 (2 measurements of 1x FEE64 carrier)
probably indicates that the spacing between the FEE64 plates is consistent. There is some variability (~0.1mm) in the Delron adaptor PCB which could affect the relative
position of the ERNI pins and sockets. Is there some variability in the width of the FEE64s (w/ rails)? |
171
|
Thu Jan 13 10:33:24 2022 |
JM, OH, TD | Adaptor Board misallignment |
The ERNI connector is not central on the adaptor boards causing misallignment with the FEE.
The adaptor board is 80mm long. The ERNI connector is 50mm long. The connector is 13mm from the 2x Lemo side and 17mm from the 3x Lemo side. This results in misallignment between the FEE and adaptor board.
To correct for misallignment, 2mm will be skimmed from one of the rails on the FEE.
Skimming the rail allows the FEE to now be inserted straight into the adaptor board connector. Attached image shows test FEE holder on the bench. FEE needs to be carefully inserted to make sure it is straight as it can now move around within the FEE holder with the 2mm skimmed off the rails. For CARME, where sight lines and space are limited, one person is required to hold and push in the FEE whilst another checks the positioning of the connectors is straight. |
Attachment 1: 20220112_160951.jpg
|
|
Attachment 2: 20220113_145105.jpg
|
|
Attachment 3: 20220113_145147.jpg
|
|
170
|
Wed Jan 12 20:07:20 2022 |
JM, OH, TD | FEE holder installation |
4 FEE holders installed on CARME. Mounting procedure devised due to limited working space to install.
Mounting procedure:
>Install FEE holder mounts (image 1) onto feedthrough flange
>Partially make up FEE holders on workbench. Only 3 cylindrical support rods, spring latches and one of the black adaptor card mounts should be attached to aid install. Bolts should be loose to aid install on CARME.
>Mount FEE holder to flange
>Connect first adaptor card into D connector on feedthrough, then screw adaptor card to black mount.
>Screw in second black adaptor card mount but do not tighten bolts. Connect second adaptor card to D connector and screw to black mount.
>Attach remaining cylindrical supports and tighten all bolts.
Bolts to install FEE holder mount onto flange (M4x25) are 2mm too long for the hole on the flange. Washer were initially used however this prevents adaptor cards being able to be secured onto thier black mounts. Bolts were shortend 2mm by davide to install. 5 shortend bolts remain in addition to thebag of unshortend M4x25 bolts.
On the nn flange of the bottom detector (non bias pin D-connector) the feedthrough is damaged. Damage may be repairable, to prevent further damage an adaptor card was not plugged into this D-connector. (image 1)
A bias test was performed for the top detector once adaptor cards are installed. First test is the bias test performed prior to adaptor cards being installed, second test was performed after installation of adaptor cards.
Installing FEEs on the outer side of the ring in the future may have some mechanical conflict with the gas target frame, see attached images. |
Attachment 1: 20220112_134113.jpg
|
|
Attachment 2: 20220112_155052.jpg
|
|
Attachment 3: 20220112_160956.jpg
|
|
Attachment 4: 20220112_161020.jpg
|
|
Attachment 5: 20220112_161034.jpg
|
|
Attachment 6: 20220112_170144.jpg
|
|
Attachment 7: 20220112_170435.jpg
|
|
Attachment 8: 20220112_165918.jpg
|
|
Attachment 9: 20220112_170048.jpg
|
|
Attachment 10: Bias2.png
|
|
168
|
Tue Jan 11 12:45:15 2022 |
TD, JM, OH | Coolant manifold installation + detector pin tests |
2x cooling water manifolds with 2x 8 ports have been installed - see attachments 1 & 2
'drain' ports at top of manifold to permit release of trapped air volumes. Local technicians will install a standard blank in one manifold and a small nipple valve on the
other so that we can bleed air in a controlled way.
#Multimeter detector pin tests.
Pins on feedthrough flanges investigated using multimeter for the top and bottom detector due to no response from bottom detector when bias was applied. Bias pins on
correct D connectors are number 78.
#Top detector
>Observe -18.5 kohm; between nn bias pin and pn bias pin (multimeter 200kohm setting)
>observe open line between nn bias pin and pins 'near' pn bias pin (multimeter 200kohm setting)
>Between random pins on same D connector (nn bias pin connector) ~0.02Mohm (multimeter 2Mohm setting)
>Between random pins and unconnected on same D connector (nn bias pin connector) - open line (multimeter 2Mohm setting)
>Between random pins and bias pin on same D connector (nn bias pin connector) ~0.01-0.02 Mohm (multimeter 2Mohm setting)
>Between nn bias pin and random pin on another D connector ~-0.4Mohm (multimeter 2Mohm setting)
#Bottom detector
>Observe open line between nn bias pin and pn bias pin (multimeter 200kohm+2Mohm; setting)
>Observe open line between nn bias pin and pins 'near' bias bias pin (multimeter 200kohm+2Mohm; setting)
>Between random pins on same D connector (nn bias pin connector) ~0.02Mohm (multimeter 2Mohm setting)
>Between random pins and unconnected on same D connector (nn bias pin connector) - open line (multimeter 2Mohm setting)
>Between random pins and bias pin on same D connector (nn bias pin connector) ~0.01-0.02 Mohm (multimeter 2Mohm setting)
>Between random pins on same D connector (pn bias pin connector) ~0.02Mohm (multimeter 2Mohm setting)
>Between random pins and unconnected on same D connector (pn bias pin connector) - open line (multimeter 2Mohm setting)
>Between random pins and bias pin on same D connector (pn bias pin connector) ~open line (multimeter 2Mohm setting)
>Between nn bias pin and random pin on another D connector ~-0.4Mohm (multimeter 2Mohm setting)
We see a similar response for both detectors except that the pn bias pin gives no response.
We have looked to find the bias pin on nearby pins to pin 78 and on the other side of the D connector (in case of cabling error-mirror or swapped cables) but cannot find.
4 FEE holders constructed and ready to install |
Attachment 1: 20220111_113628.jpg
|
|
Attachment 2: 20220111_113613.jpg
|
|
Attachment 3: 20220111_111428.jpg
|
|
167
|
Mon Jan 10 20:52:53 2022 |
JM, TD, OH | Baking tent removed, VI tests |
Baking tent, frame and trays fully removed from CARME chamber. Removal of baking frame required removing air lines from upper gate valve and repositioning of vacuum hose connecting turbo pump to backing line and turbo power cable. Some difficulty encountered removing upper baking frame now the target support frame is assembled. Small modification seperating section in two would aid future installation and disammebly.
Baking trays removed with difficulty. Main baking tray bent to extract from the frame -> updated design required for future bakeout cycle.
VI tests performed on the upper and lower detectors. Upper detector shows characteristic curve (attached plot), lower detector shows no response to increased voltage. Multimeter used on adapter boards with resistance of ~18.5 kΩ observed between bias pins for the upper detector. Open line seen between bias pins on the lower detector, some connections observed for other pins on the board -> requires further investigation.
NOTE - Channel 0 on bias supply is unstable, should not be used.
Most equipment for installation of FEE's and coolant manifolds found, assembley of FEE holders started. Full list of current swagelock inventory as follows:
SS-QTM2-D-8M0 MALE QUICK DISCONNECT |
15 |
SS-8M0-1-6RS MNIFOLD ADAPTOR |
40 |
NY-8M4-1 ferrule fronts (nylon) |
120 |
NY-8M3-1 ferrule rears (nylon) |
120 |
SS-QTM2A-B-8M0 FEMALE QUICK DISCONNECT |
15 |
|
Attachment 1: 20220110_175547.jpg
|
|
Attachment 2: 20220110_175811.jpg
|
|
Attachment 3: vi-curve.png
|
|
166
|
Tue Dec 7 09:39:59 2021 |
CB (remotely) | Gate valves closed |
Dry run complete.
YR09 section valves closed. Section isolated. Manual valve between electron target and CARME remain open. |
165
|
Tue Nov 30 10:49:43 2021 |
CB (remotely) GV | Gate valve opened |
GSI vacuum division IE514 gauge connected.
Gleb reports gate downstream opened. Vacuum goes up and recovers as expected. |
Attachment 1: vacplot24.png
|
|
164
|
Sun Nov 28 15:57:45 2021 |
CB (remotely) | Nov 2021 temperatures |
Attached |
Attachment 1: 211128164424_temperatures.zip
|
163
|
Fri Nov 26 10:01:45 2021 |
CB (remotely) ML | Power cut in the CRYRING cave |
ML reports power cut in CRYRING cave. Reasons unknown. No effect on vacuum, but top valve closed for safety reason. Ramping down temperature. |
162
|
Wed Nov 24 12:08:08 2021 |
CB + ML OG | Ion pumps on |
Found SIP power unit for D2000 #4 not functioning correctly. Max current regardless of ion element connected to.
ML kindly loaned a spare. All D2000 work Ok now. Will have to send faulty unit for repairs back to SAES.
Opened up front part of tent just enough to allow GSI vacuum division to connect IE 514 and GSI ion pump. Tent to remain until January.
Left to do this year:
- Turn on GSI ion pump
- Turn on GSI IE514
- Close top gate valve
- Turn off maglev turbo (very close to magnet) |
Attachment 1: vacplot24.png
|
|
Attachment 2: vacplot.png
|
|
Attachment 3: CARMEtemps.png
|
|
161
|
Tue Nov 23 08:49:33 2021 |
CB + ML | NEG activation completed |
Found temperatures and pressure stabilised overnight. See attached.
12:00 Began activation of 3x UHV1400 in channel 2, top module. Started using 90 W as per procedure. No effect on temperature. Increased to 120 W. OK
14:30 Began activation of 2x UHV1400 in channel 1, in front of the detectors. Considered moving detectors but decided against it since bellows are rated to move at 20C at most. Activated at 60 W (vs. 75 W in procedure). OK
All NEG elements off. CARME heaters ramping down to 60 C overnight.
Ion pumps activated < 100 C. Valve closed. #4 doesn't come online properly. Sone issues with #2 as well but it recovered on its own. Left cooling down. No changes. Ion pumps off and valve open. Will try tomorrow. Could be faulty cable?
Vacuum division may want to activate their pump as well - times unclear. |
Attachment 1: vacplot24.png
|
|
Attachment 2: CARMEtemps.png
|
|
Attachment 3: vacplot24-1.png
|
|
Attachment 4: CARMEtemps-1.png
|
|
160
|
Mon Nov 22 08:06:01 2021 |
RSS CB + ML | NEG activation starts |
08:20 Temperature of channels 1-3 and 5-6 were reduced to 120 degrees. Immediately after, the conditioning of the NEGs was stopped and D2000s ion pumps were flashed 5 times. The temperature of the gas jet target and the electron target channels was reduced to 120 degrees.
11:00 Started activation of D2000 following procedure in ... at T=400 C for trise 30 min + thold 90 minute. After activation, NEG elements are left at 200 C to maintain cleanliness. Temperatures increasing, but still well below safety margins.
16:00 Started activation of NEGs from channel 1 at 200 W. No effects on temperatures. Activated channel 2 as well. No significant effects. Increased activation power to 250 W. No issues
16:30 Given temperatures are still low decided to activate 3 and 4 with 230 W (Vs. 150 W last time). Ok for 30 minutes then temperature starts increasing a bit too quickly. Decreased to 200 W. Ok.
18:45 Activation all NEGs on middle module complete. Left overnight to stabilise / cool.
21:00 Channel 6 of the heater stops responding properly. ML asked HKR to disable. No effects on temperatures.
|
Attachment 1: IMG_5674.JPG
|
|
Attachment 2: CARMEtemps.png
|
|
Attachment 3: vacplot.png
|
|